# Artificial Intelligence in Image-based Cardiovascular Disease Analysis: A Comprehensive Survey and Future Outlook

### v3

@@ 3.3.2 Coronary Arteries
Coronary Flow Reserve (CFR) in Cardiac Assessment. CFR is an essential measure in cardiology, gauging the coronary arteries’ capacity to augment blood flow in response to increased cardiac demand, as depicted in Figure 4 (d). This metric is particularly useful in evaluating patients who exhibit ischemic symptoms without significant stenosis, helping to guide decisions on revascularization or medication management. CFR’s estimation can be achieved non-invasively through Transthoracic Doppler Echocardiography, MRI, and PET, offering a safer alternative for certain patient groups [194]. Unlike Fractional Flow Reserve (FFR) that evaluates the significance of arterial narrowing, and the Index of Microcirculatory Resistance (IMR) that measures microcirculation resistance, CFR provides an overarching view of the entire coronary circulatory system’s flow capacity, encompassing both large and small vessels. Typically, CFR is assessed through invasive coronary catheterization, where blood flow velocity is measured at rest and during induced hyperemia using specialized Doppler guidewires or pressure sensors.
Fractional Flow Reserve (FFR) in Coronary Assessment. FFR is an index for quantifying blockage severity in coronary arteries, typically determined through invasive catheter-based methods, as illustrated in Figure 1 (e). However, modern approaches are exploring non-invasive FFR estimation, using anatomical data from CT scans of the heart and coronary arteries. Despite their promise, these methods, mainly physics-based models, face challenges due to high computational demands, limiting their routine clinical application.
@@
With the integration of computational AI, there’s a growing interest in developing non-invasive FFR estimation techniques from medical imaging modalities like coronary computed tomography angiography (CCTA). These AI-driven methodologies are aimed at circumventing the need for invasive procedures, making CAD assessment more patient-friendly and efficient. As seen in Figure 4 (c, d), a typical AI-powered FFR analysis workflow includes segmentation of coronary arteries within CTA images to delineate their anatomical structure [195, 196, 197]. This process is often supplemented with automated branch labeling systems [198], enhancing lesion localization accuracy in coronary arteries [199, 84, 200, 201]. The culminating step involves leveraging deep learning models, trained on the segmented and labeled structural and functional features, to predict FFR. These AI models present a compelling alternative to traditional physics-based methods, offering potentially quicker and equally accurate FFR estimations [202, 203].
@@
Coronary Microvascular Disease (CMVD). CMVD disease within the coronary circulatory system, particularly affecting the microvasculature, can lead to clinical events in patients without obvious epicardial coronary stenosis [204]. CMVD specifically targets the heart’s small blood vessels or microcirculation, often resulting in symptoms like chest pain and shortness of breath, though it can also be asymptomatic. Unlike obstructive coronary artery disease, which impacts larger vessels, CMVD involves reduced blood flow through these microvessels (See Figure 4 (d)). Traditional angiography typically fails to detect CMVD, as it mainly visualizes larger coronary arteries, necessitating alternative diagnostic approaches.
Index of Microcirculatory Resistance (IMR) in Microvascular Assessment. IMR serves as a key invasive metric for directly evaluating coronary microvascular function. This index is obtained using specialized pressure and temperature-sensitive wires, similar to those used in FFR measurements. It’s crucial to recognize that high FFR values, typically above 0.80, do not entirely rule out the risk of future clinical events. Patients with high FFR readings may still experience clinical issues, emphasizing the need for comprehensive assessments beyond FFR [23].
IMR is particularly insightful for diagnosing microvascular dysfunction, proving valuable in both stable patients and those with acute or recent MI [205]. Despite its utility, the medical field still seeks reliable and non-invasive quantitative techniques for diagnosing CMVD. Developing such methods is essential for selecting appropriate treatments for coronary artery disease and improving coronary microcirculation outcomes [206]. Innovation in this area includes the recent introduction of a novel parameter, ’caIMR,’ aimed at characterizing the physiological effects of microcirculatory dysfunction [207]. This novel metric potentially offers critical insights into microvascular diseases, enhancing patient evaluation and treatment strategies. The application of ’caIMR’ could represent a significant advancement in understanding and managing microvascular conditions.
3.4 Connections between Non-vessel and Vessel Structures
The heart’s functionality is intricately dependent on both its non-vessel and vessel structures. Understanding the interplay between these two types of structures is pivotal for a holistic comprehension of cardiac function, specifically how structural integrity and blood supply interact.
As illustrated in Figure 5 (Top), the myocardium’s dependency on a continuous supply of oxygen and nutrients, primarily delivered by the coronary arteries, is a key aspect of cardiac function. Disruptions in this supply, such as from atherosclerosis or thrombosis in the coronary arteries, can result in myocardial ischemia or infarction, impairing the heart’s pumping efficiency [3]. Conversely, changes in non-vessel structures, like ventricular hypertrophy or valvular diseases, can influence the heart’s blood demand and workload, subsequently affecting coronary circulation. For instance, ventricular hypertrophy heightens the myocardium’s oxygen needs, which can lead to an imbalance if the coronary arteries are unable to meet this increased demand.
There exists a dynamic feedback loop between myocardial performance and coronary vessel health. For example, myocardial damage from infarction can alter blood flow dynamics, impacting cardiac output and the stress on coronary vessels [179]. The cardiac cycle involves myocardial contraction and relaxation, as well as valve operation, all of which are vital for effective blood propulsion through the coronary vessels and into systemic circulation. Structural or functional abnormalities in these non-vessel components can hamper blood circulation efficiency, thereby affecting overall cardiac function [64]. As shown in Figure 5 (Bottom), the myocardial computed tomographic perfusion (CTP) derived fractional flow reserve (CTP-FFR) models have been initially established [182, 181]. The models demonstrate superior diagnostic capabilities for coronary artery disease (CAD), particularly in cases of mild to moderate stenosis.
A recent work developed a multiscale, patient-specific model to enable blood flow simulation from major coronary arteries down to myocardial tissue [208]. Specifically, a stand-alone coronary model and an integrated coronary-myocardium coupled model were developed and examined with the objective of simulating myocardial perfusion under both healthy and pathological conditions. The patient’s vasculature is initially segmented from coronary computed tomography angiography data and then extrapolated to the arteriole level using a network of synthetic trees. Then, the CT-FFR blood flow is modeled by integrating a coronary artery model with a single-compartment Darcy myocardium model for the joint hyperemic Myocardial Blood Flow (MBF) for the coupled model. Finally, the predicted and true perfusion maps are compared. This work represents an instance of a computational model simulating blood flow from epicardial coronary arteries to the left ventricle myocardium, applied and validated using human data.
Currently, the application of AI in comprehensively understanding the connection between non-vessel and vessel structures remains limited. However, such understanding is critical for diagnosing and treating cardiac diseases, as it aids in identifying the root causes of cardiac dysfunction, informs surgical intervention planning, and supports long-term cardiac health management strategies.
Figure 6: Left: Cardiac Imaging Genetic Analysis Pipeline: This diagram illustrates the process of extracting detailed features of heart structure and function using AI models from cardiac imaging data, followed by the exploration of endophenotypic and genetic associations through Genome-wide association studies (GWAS), with a consideration of lifestyle and environmental factors. Right: Overview of the study, which used Cardiovascular magnetic resonance (CMR) and brain MRI traits as endophenotypes to explore the phenotypic and genetic connections between the heart and the brain [16].
@@ 4 Population Image-based CVD Analysis
Our exploration in this section delves into understanding cardiovascular diseases (CVDs) through detailed population-based imaging studies, utilizing AI and statistical tools. We address the diversity of CVDs, encompassing various disorders like coronary artery disease, heart failure, and arrhythmias, each with its own clinical and pathophysiological traits. Acknowledging the significant influence of lifestyle and behavioral factors on these diseases, our focus extends to how these elements, along with genetics and age, shape CVD risks [209]. Population-based study is important to construct temporal Causal Genetic-Imaging-Clinical (CGIC) pathways to examine the causal relationships between genetic predispositions, cardiac imaging features, and clinical outcomes in CVDs. This integrative approach is designed to deepen our understanding of CVD etiology and to enhance strategies for prevention, diagnosis, and treatment.
@@ 4.1 Cardiac Imaging Genetics
The intricate genetic architecture of CVDs has been elucidated through numerous genome-wide association studies (GWAS) over the past decade and a half, highlighting the interplay of genetic and environmental factors in these diseases. Studies cataloged in the GWAS Catalog [210] reveals that nearly 1,700 studies on 450 CVD-related phenotypes have been conducted. This research has advanced our understanding of cardiovascular biology and pathophysiology, uncovering complex interactions between genetic and non-genetic risk factors. Overall, these studies underline the importance of continued exploration into the genetic basis of CVDs to enhance prevention, diagnosis, and treatment strategies.
@@
Case-control GWAS is the most direct method to discover common genetic risk factors underlying disease. GWAS initially focused on common complex diseases like Coronary Artery Disease (CAD) and myocardial infarction (MI), atrial fibrillation (AF), heart failure, stroke, and hypertension. For instance, although the heritability of CAD/MI is estimated to be around 40-50%, only about 30% of this heritability is explained by the currently identified Single Nucleotide Polymorphisms (SNPs) [212, 213]. This suggests that there are still genetic factors of CAD/MI yet to be discovered.
For more rare cardiovascular conditions, GWAS have begun to reveal the role of common genetic variants as either modifiers of disease penetrance and severity in cases with rare Mendelian variants or as substantial contributors to overall disease risk in non-Mendelian cases [214, 215]. This highlights the need for further research and collaboration in understanding these diseases.
The development of Polygenic Risk Scores (PRS), which summarize the aggregate effect of all identified risk variants for a disease, has been a notable advancement. PRS have shown potential for risk stratification, especially when combined with traditional risk factors [216]. However, the integration of PRS into clinical practice requires further research to understand its benefits and workflows.
Imaging genetics in CVDs integrates advanced imaging and genetic analysis to decode the genetic underpinnings of heart conditions [217, 218]. This field examines the impact of genetic variations on cardiovascular traits observed through medical imaging, such as MRI or CT scans, facilitating the identification of genetic markers linked to CVD development and progression. An overview of the heart imaging genomics pipeline is shown in Figure 6 (left). This method proves invaluable for complex diseases like heart failure, coronary artery disease, and arrhythmias, aiding in early risk detection, understanding disease mechanisms, and tailoring treatment strategies to individual genetic profiles.
The inclusion of cardiac structural and functional measures in CGIC pathways is justified by three key factors. Firstly, genetic variations more significantly influence cardiac imaging traits than the broader, more heterogeneous categories of behaviorally defined diseases (e.g., Ischemic heart disease, heart failure, and hypertensive heart disease). Secondly, understanding the genetic variations related to cardiac changes enhances our grasp of the mechanisms that predispose individuals to such behaviorally defined disorders. Thirdly, the difficulty in gathering sufficiently large cohorts for rare diseases makes the study of disease-relevant endophenotypes in large biobank cohorts essential for discovering genetic influences on these rare conditions.
The enhanced power of ECG trait GWAS to detect substantially more genetic associations compared with rare disease case-control studies may provide more informative PRS, new insights into disease mechanisms and pathways, and the potential to identify novel Mendelian disease genes or drug targets.
Recent GWAS have analyzed numerous ECG quantitative traits across tens of thousands of individuals, leading to the identification of hundreds of significant genetic associations [219, 220]. These studies have focused on traits crucial for cardiac functions, such as heart rate, PR interval, QRS duration for cardiac conduction, and QT interval for repolarization. The significant findings in these GWAS have underscored the importance of endophenotype GWAS in understanding cardiac diseases, especially when strong signals are identified at genes known to be associated with arrhythmias, suggesting these may be the causal genes. For instance, variations at key Long QT Syndrome (LQTS) genes like KCNQ1, KCNH2, and SCN5A have shown strong associations with the QT interval, and SCN5A with PR and QRS traits as well, linking these to both common and rare cardiac conduction diseases. The correlation between the effect sizes for QT interval-associated Single Nucleotide Polymorphisms (SNPs) in the general population and disease risk in LQTS highlights the relevance of such studies. The ability of ECG trait GWAS to uncover more genetic associations compared to rare disease case-control studies may enhance PRS, provide new insights into disease mechanisms, and potentially identify novel Mendelian disease genes or drug targets.
In cardiomyopathies, the key endophenotypes involve traits that delineate the heart’s structure and function, best assessed via advanced imaging as discussed in Sections 2 and 3. The recent availability of large-scale datasets (e.g., the UK Biobank (UKB)) marks significant progress, despite the complexity and cost associated with these techniques. The critical endophenotypes for Hypertrophic Cardiomyopathy (HCM) and Dilated Cardiomyopathy (DCM) concern the left ventricle (LV), including its mass and wall thickness, myocardial deformation, volumes, and functional measures such as ejection fraction. Compared with traditional cardiomyopathy case-control studies, initial imaging genetics analyses have uncovered numerous genome-wide significant associations [221, 222]. It’s notable that many loci are linked with multiple LV traits, underscoring their genetic and phenotypic correlation. Importantly, some loci are proximal to established Mendelian cardiomyopathy genes, like TTN and PLN, reinforcing the endophenotypes’ relevance in uncovering biological processes common to cardiomyopathies. PRS based on LV endophenotypes have shown predictive value for heart failure and DCM in UKB subsets.
Pirruccello et al.’s [223] pioneering study leveraged deep learning to analyze left atrial characteristics from cardiovascular MRI scans of over 40,000 UKB participants, focusing on key metrics such as minimum and maximum volumes, stroke volume, and emptying fraction. Their subsequent GWAS on individuals without prior cardiovascular conditions revealed 20 genetic loci associated with LA traits, illuminating the genetic underpinnings of LA structure, function, and atrial fibrillation risk. Another study by the same group extended this methodology to the right heart chamber, identifying and measuring the right atrium, ventricle, and pulmonary artery [224]. This comprehensive analysis identified 130 genetic loci associated with right heart measurements, some near genes linked to congenital heart disease, and highlighted the potential of genome-wide polygenic predictors in understanding the genetic basis of right heart structure and function, offering new insights into cardiovascular diseases and their treatment.
A recent study [225] introduced groundbreaking findings on the relationship between CMR measured Pericardial Adipose Tissue (PAT) and various cardiovascular structure and function indicators. The study analyzed PAT data from over 42,000 UKB participants. It delved into the connections between PAT and the risk factors for heart failure, ischemic heart disease (IHD), adverse LV adiposity, and AF. Significantly, this research sought to ascertain whether these associations remain distinct from conventional vascular risk factors and different obesity measures, including body metrics, impedance-based fat assessments, and MRI evaluations of abdominal visceral fat.
Dataset Target Task Year Modality Disease
LAScarQS [226] LA, LA Scar Segmentation 2022 LGE MRI Atrial fibrillation, etc
CMRxMotion [227] LV, Myo Segmentation 2022 MRI Respiratory Motion, etc
MyoPS 2020 [228] LV, RV, Myo Segmentation 2020 Cine MRI Myocardial Infarction, etc
MSD Cardiac [229] LA Segmentation 2019 MRI CVD, Atrial fibrillation, etc
MM - Whole Heart Segmentation [230] Whole Heart Segmentation 2019 MRI, CT Atrial fibrillation, etc
CAMUS [24] LV LA Segmentation 2019 2D Ultrasound ejection fraction
Atria Segmentation challenge [231] LA Segmentation 2018 LEG MRI Atrial fibrillation
Left Ventricle Full Quantification Challenge [232] LA Quantification 2018 Cine MRI General CVD
ACDC, MICCAI challenge [233] LV, RV, Myo Segmentation 2017 Cine MRI Mycardial infarction, etc
HVSMR, MICCAI workshop [234] Blood pool, Myo Segmentation 2016 3D MRI Congenital heart defects
CETUS, MICCAI challenge [235] LV Segmentation 2014 3D Ultrasound Myocardial infarction
Left Atrial Segmentation challenge [236] Left Atrial Segmentation 2013 3D MRI General CVD
Left Ventricle Infarct Segmentation [237] Left Ventricle Segmentation 2012 LGE MRI Ischemic cardiomyopathy
CDEMRIS fibrosis scar challenge [238] LA fibrosis and scar Segmentation 2012 LGE MRI Atrial fibrillation
Right Ventricle Segmentation challenge [239] RV Segmentation 2012 Cine MRI Myocarditis, etc
LV Segmentation Challenge (LVSC) [240] LV, Myo Segmentation 2011 Cine MRI Coronary artery disease, etc
Cardiac Atlas Project (CAP) [241] Whole Heart Multi-tasks 2011 MRI General CVD
CardioVascular Research Grid (CVRG) [242] Whole Heart Multi-tasks 2011 Multi-modality General CVD
Sunnybrook Cardiac Data [243] LV, Myo Segmentation 2009 Cine MRI Hypertrophy, heart failure, etc
York [244] LV, Myo Segmentation 2008 Cine MRI Cardiomyopathy, etc
Aortic Stiffness 4D Flow (AS4DF) [245] Aorta Hemodynamics 2021 MRI Arterial Hypertension, etc
ImageCAS [197] Coronary artery Segmentation 2022 CTA Vessel stenosis, etc
CLS12 [104] Coronary artery Segmentation 2012 CTA Stenosis Analysis
Coronary Artery Tracking Challenge [246] Coronary artery Segmentation 2008 CTA Calcium Analysis
TABLE I: Summary of public cardiac image datasets on cardiac image analysis. Top: Non-vessel structures and Bottom: Vessel structures.
A new study [247] embarked on uncovering the genetic factors that influence the dimensions of key cardiovascular structures, including the left ventricular outflow tract (LVOT), aortic root, and ascending aorta, noted for their complex structure and unique developmental origins. The study meticulously measured diameters at specific points within these structures to conduct GWAS aimed at identifying associated genetic variants. Furthermore, the study by Aung [248] delved into the genetic basis of RV structure and function. This research, involving 29,506 UKB participants and leveraging cardiovascular MRIs analyzed by advanced deep learning algorithms, identified 25 unique genetic loci associated with RV phenotypes. Notably, some candidate genes were found to overlap with genes linked to Mendelian cardiomyopathy, shedding light on their roles in cardiac muscle contraction and cellular adhesion. Additionally, the study highlighted the association of RV PRS with DCM and CAD, offering new insights into the genetic foundations of RV-related cardiac conditions.
In summary, utilizing AI models allows for exploring intricate relationships between genetic variations and imaging characteristics, facilitating the identification of new associations and the creation of improved predictive models. AI algorithms can effectively amalgamate genetic and imaging data, enabling the delineation of distinct disease subtypes grounded in both genetic profiles and imaging attributes [249].

### v4

@@ 3.4.2.2 Coronary Artery Diseases (CAD)
Coronary Flow Reserve (CFR) is an essential measure in cardiology, gauging the coronary arteries’ capacity to augment blood flow in response to increased cardiac demand, as depicted in Figure 3 (d). This metric provides an overarching view of the entire coronary circulatory system’s flow capacity, encompassing both large and small vessels. It is particularly useful in evaluating patients who exhibit ischemic symptoms without significant stenosis, helping to guide decisions on revascularization or medication management. CFR’s estimation can be achieved non-invasively through MRI and PET, etc, offering a safer alternative for certain patient groups [fawaz2023invasive]. Typically, CFR is assessed through invasive coronary catheterization, where blood flow velocity is measured at rest and during induced hyperemia using specialized Doppler guidewires or pressure sensors.
@@
Fractional Flow Reserve (FFR) is an index for quantifying blockage severity in coronary arteries, typically determined through invasive catheter-based methods, as illustrated in Figure 1 (e). However, modern approaches are exploring non-invasive FFR estimation, using anatomical data from CT scans of the heart and coronary arteries. Despite their promise, these methods, mainly physics-based models, face challenges due to high computational demands, limiting their routine clinical application. With the integration of computational AI, there’s a growing interest in developing non-invasive FFR estimation techniques from medical imaging modalities like coronary computed tomography angiography (CCTA). These AI-driven methodologies are aimed at circumventing the need for invasive procedures, making CAD assessment more patient-friendly and efficient. As seen in Figure 3 (c, d), a typical AI-powered FFR analysis workflow includes segmentation of coronary arteries within CTA images to delineate their anatomical structure [zheng2011machine, kong2020learning, zeng2022imagecas]. This process is often supplemented with automated branch labeling systems [li2023automated], enhancing lesion localization accuracy in coronary arteries [wu2019automated, yang2020cpr, zhao2023agmn, li2020segmentation]. The culminating step involves leveraging deep learning models, trained on the segmented and labeled structural and functional features, to predict FFR. These AI models present a compelling alternative to traditional physics-based methods, offering potentially quicker and equally accurate FFR estimations [itu2016machine, gao2020learning].
3.4.2.3 Coronary Microvascular Disease (CMVD)
@@
CMVD disease within the coronary circulatory system, particularly affecting the microvasculature, can lead to clinical events in patients without obvious epicardial coronary stenosis [ayub2020coronary]. CMVD specifically targets the heart’s small blood vessels or microcirculation, often resulting in symptoms like chest pain and shortness of breath, though it can also be asymptomatic. Unlike obstructive coronary artery disease, which impacts larger vessels, CMVD involves reduced blood flow through these microvessels (See Figure 3 (d)). Traditional angiography typically fails to detect CMVD, as it mainly visualizes larger coronary arteries, necessitating alternative diagnostic approaches. The Index of Microcirculatory Resistance (IMR) serves as a key invasive metric for directly evaluating coronary microvascular function. This index is obtained using specialized pressure and temperature-sensitive wires, similar to those used in FFR measurements. It’s crucial to recognize that high FFR values, typically above 0.80, do not entirely rule out the risk of future clinical events. Patients with high FFR readings may still experience clinical issues, emphasizing the need for comprehensive assessments beyond FFR [berry2014fractional]. IMR is particularly insightful for diagnosing microvascular dysfunction, proving valuable in both stable patients and those with acute or recent MI [zhao2023early]. Developing such methods is essential for selecting appropriate treatments for coronary artery disease and improving coronary microcirculation outcomes [gao2022clinical].
3.5 Connections of Non-vessel and Vessel Structures
The heart’s functionality is intricately dependent on both its non-vessel and vessel structures. Understanding the interplay between these two types of structures is pivotal for a holistic comprehension of cardiac function, specifically how structural integrity and blood supply interact. The myocardium’s dependency on a continuous supply of oxygen and nutrients, primarily delivered by the coronary arteries, is a key aspect of cardiac function. Disruptions in this supply, such as from atherosclerosis or thrombosis in the coronary arteries, can result in myocardial ischemia or infarction, impairing the heart’s pumping efficiency [li2023multi]. There exists a dynamic feedback loop between myocardial performance and coronary vessel health. For example, myocardial damage from infarction can alter blood flow dynamics, impacting cardiac output and the stress on coronary vessels [fukushima2020patient]. Structural or functional abnormalities in these non-vessel components can hamper blood circulation efficiency, thereby affecting overall cardiac function [slomka2021quantitative]. A recent work developed a multiscale, patient-specific model to enable blood flow simulation from major coronary arteries down to myocardial tissue [papamanolis2021myocardial]. Specifically, a stand-alone coronary model and an integrated coronary-myocardium coupled model were developed and examined with the objective of simulating myocardial perfusion under both healthy and pathological conditions. Then, the CT-FFR blood flow is modeled by integrating a coronary artery model with a single-compartment Darcy myocardium model for the joint hyperemic Myocardial Blood Flow (MBF) for the coupled model. Finally, the predicted and true perfusion maps are compared. This work represents an instance of a computational model simulating blood flow from epicardial coronary arteries to the left ventricle myocardium, applied and validated using human data. Currently, the application of AI in comprehensively understanding the connection between non-vessel and vessel structures remains limited. However, such understanding is critical for diagnosing and treating cardiac diseases, as it aids in identifying the root causes of cardiac dysfunction, informs surgical intervention planning, and supports long-term cardiac health management strategies.
3.6 Section Summary
Section 3 reviews how modern AI methods are applied across the cardiac imaging pipeline for both non-vessel structures (atria, ventricles, myocardium) and vessel structures (aorta, coronary arteries, microvasculature. AI supports key tasks such as segmentation, motion tracking, feature extraction, stenosis detection, and physiologic assessment (e.g., FFR, CFR, IMR). It also highlights the need for AI models that jointly consider structural and functional information, reflecting the physiological coupling between myocardium and coronary circulation. Overall, Section 3 shows how AI can enhance diagnostic accuracy and clinical decision-making across the spectrum of cardiovascular disease.
@@ 4 Population Image-based CVD Analysis
Findings from population-based imaging studies provide a foundation for individual analysis by identifying broader risk factors and disease progression patterns, which can be refined for personal application. The connection between individual and population-based studies is in their shared goal of improving cardiovascular health, but they complement each other by bridging individual and group insights. Conversely, individual analysis can inform population studies by revealing which subgroups might benefit most from certain interventions, further refining predictive models. Together, these approaches enhance our understanding of CVD and support the transition from broad-spectrum therapies to personalized, data-driven care that considers both individual and population health factors. Thus, our exploration in this section delves into understanding CVDs through detailed population-based imaging studies, utilizing AI and statistical tools. We address the diversity of CVDs, like coronary artery disease, heart failure, and arrhythmias, each with its own clinical and pathophysiological traits. Acknowledging the significant influence of lifestyle and behavioral factors on these diseases, our focus extends to how these elements, along with genetics and age, shape CVD risks [sing2003genes]. Population-based study is important to construct temporal Causal Genetic-Imaging-Clinical (CGIC) pathways to examine the causal relationships between genetic predispositions, cardiac imaging features, and clinical outcomes in CVDs [zhao2023heart]. This integrative approach is designed to deepen our understanding of CVD etiology and to enhance strategies for prevention, diagnosis, and treatment.